A 80-kg man jumps out a window into a fire net 10 m below. The net stretches 2.0 m before bringing him to rest and tossing him back into the air. What is the potential energy of the stretched net if nonconservative forces dissipate no energy?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Mass of man = 80 kg
Hence total weight. = m *g = 80 *9.8 = 784 N
 
The man's energy of falling is arrested when he has fallen 10 meters to the net plus another 2 meters of stretching the net.  His weight is 784 N, so the total energy is: 

784 N * (10 m + 2 m) 
= 9408 N-m = 9408 J 

If the net is fully elastic, the potential energy of the net is equal to the gravitational energy absorbed.